OpenStudy (anonymous):

Can someone please simplify sqrt.180-sqrt.20? Thank you.

OpenStudy (anonymous):

4sqrt5

OpenStudy (anonymous):

Can you show me the steps please?

OpenStudy (anonymous):

You just have to find the largest sq rt number that goes into 180 which is 36. The sq rt of 36 is 6. Then same thing but with 20! So 4 is the largest, and sq. root is 2

jimthompson5910 (jim_thompson5910):

The key is to factor each radicand where one number is a perfect square. Then you break up the roots and simplify like this: \[\Large \sqrt{180} - \sqrt{20}\] \[\Large \sqrt{36*5} - \sqrt{4*5}\] \[\Large \sqrt{36}*\sqrt{5} - \sqrt{4}*\sqrt{5}\] \[\Large 6*\sqrt{5} - 2*\sqrt{5}\] \[\Large (6 - 2)*\sqrt{5}\] \[\Large 4*\sqrt{5}\]

OpenStudy (anonymous):

Ok, we got it now. Thank you all for all of the help!